### v1

# Is Transductive Learning Equivalent to PAC Learning?

Abstract
Most work in the area of learning theory has focused on designing effective Probably Approximately Correct (PAC) learners. Recently, other models of learning such as transductive error have seen more scrutiny. We move toward showing that these problems are equivalent by reducing agnostic learning with a PAC guarantee to agnostic learning with a transductive guarantee by adding a small number of samples to the dataset. We first rederive the result of Aden-Ali et al. [1] reducing PAC learning to transductive learning in the realizable setting using simpler techniques and at more generality as background for our main positive result. Our agnostic transductive to PAC conversion technique extends the aforementioned argument to the agnostic case, showing that an agnostic transductive learner can be efficiently converted to an agnostic PAC learner. Finally, we characterize the performance of the agnostic one inclusion graph algorithm of Asilis et al. [4] for binary classification, and show that plugging it into our reduction leads to an agnostic PAC learner that is essentially optimal. Our results imply that transductive and PAC learning are essentially equivalent for supervised learning with pseudometric losses in the realizable setting, and for binary classification in the agnostic setting. We conjecture this is true more generally for the agnostic setting.
1 Introduction
The dominant paradigm in statistical learning theory is the probably approximately correct (PAC) model due to Valiant [32]. On the other hand, the transductive model of learning, as originally employed by Haussler et al [20], does away with distributional assumptions and evaluates a learner’s leave-one-out performance on an adversarially-chosen sample. The elegant simplicity of the transductive model has enabled application of combinatorial and graph-theoretic insights to learning, as evidenced by a rich body of work which employs transductive learners as precursors to PAC learners (e.g. [20, 13, 1]). Indeed, some of this recent literature suggests that transductive and PAC learning might be equivalent in a rich variety of learning settings, with essentially identical sample complexities.
The present paper examines the strength of this equivalence, not only in the poorly understood agnostic setting, but also at a greater level of generality in the realizable setting. We build on the recent work of Aden-Ali et al. [1] to show that both realizable and agnostic PAC learning can be reduced to transductive learning so long as the loss function is a pseudometric. This reduction comes at modest expense: a small number of additional samples determined by the error and confidence parameters, as well as multiplication of the error by a universal constant. This implies a strong sense of equivalence of the two models in the realizable setting, where it is known, and easy to show, that transductive learning reduces to PAC learning at the cost of a constant factor in the error. Such an essentially loss-less reduction from transductive to PAC learning is not known in the agnostic setting, where it remains plausible that transductive learning is fundamentally more difficult than its PAC counterpart. [^1] Indeed, the naive reduction from transductive to PAC learning loses a factor of $1/\epsilon$ in the sample complexity, where $\epsilon$ is the required error rate. We rule out such a separation for binary classification by explicitly quantifying the performance of the optimal agnostic transductive learner in terms of the VC dimension. This implies that transductive learning serves as a precursor to an essentially optimal PAC learner in agnostic binary classification, via our reduction. We view this as justifying a conjecture that transductive and PAC learning are equivalent more broadly in the agnostic setting.
1.1 Background
The overarching goal of statistical learning theory is to “learn” structure from data in a manner that generalizes to unseen future data. In supervised learning, there is a data domain $\mathcal{X}$ and label set $\mathcal{Y}$. The learner observes a sequence of labeled data points $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in(\mathcal{X}\times\mathcal{Y})^{n}$, called the sample or training data, from which it derives a predictor $h:\mathcal{X}\to\mathcal{Y}$ which anticipates the label of a future test data point. The quality of the prediction is evaluated using a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, applied to the predicted label and the true (unseen) label. The loss of a predictor is compared, in some aggregate sense, to that of the best predictor in some benchmark class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ of hypotheses.
Since the inception of learning theory, several different data models and benchmarks have been used to evaluate learners [32, 18, 9, 33, 24, 11]. The most prominent theoretical framework for supervised learning is the PAC (Probably Approximately Correct) model due to Valiant [32]. In addition to the hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, a learning problem in this setting fixes a class of plausible data distributions supported on $\mathcal{X}\times\mathcal{Y}$. An instance of the learning problem is then determined by an (unknown) distribution $\mathcal{D}$ in this class, from which both training and test data are drawn. When all data distributions are plausible, this is the agnostic setting of learning. When only distributions that are consistent with some hypothesis are plausible, meaning that there exists $h\in\mathcal{H}$ such that $(x,y)\sim\mathcal{D}$ satisfies $y=h(x)$ with probability $1$, this is the realizable setting. For functions $\epsilon,\delta:\mathbb{N}\to\mathbb{R}_{\geq 0}$, a learner is said to be an $(\epsilon,\delta)$-PAC learner if, given training data consisting of $n$ i.i.d. draws from $\mathcal{D}$, with probability at least $1-\delta(n)$ it outputs a predictor whose expected loss on test data from $\mathcal{D}$ is at most $\epsilon(n)$ in excess of that of the best predictor in $\mathcal{H}$. Inversely, the sample complexity $m_{\text{PAC}}(\epsilon,\delta)$ of a PAC learner quantifies the minimum number of samples needed to guarantee excess loss $\epsilon$ with probability $1-\delta$, for given $\epsilon,\delta\geq 0$.
Extensive research has been conducted on the sample complexity function for various supervised learning problems, whose existence defines the PAC learnability of a problem. The primary focus of PAC learnability theory is to characterize the problem instances that can be learned in this framework across various domains, such as binary classification [19, 10], multi-class classification [25, 26, 13, 12, 14], regression [30, 8, 14, 6], and beyond [5]. In addition to these characterizations, a rich literature has examined upper bounds [15, 18, 9, 8, 14, 6] and lower bounds [32, 30] on the sample complexity of PAC learning problems.
In the transductive model of learning there is also a hypothesis class $\mathcal{H}$, but no distribution. An instance of the problem consists of a sample $S\in(\mathcal{X}\times\mathcal{Y})^{n}$ that is selected by an adversary. Of the $n$ labeled data points in this sample, $n-1$ are selected uniformly at random as the training set, and a predictor is applied to the remaining test point. The realizable setting of transductive learning restricts attention to samples that are consistent with some hypothesis in $\mathcal{H}$. The agnostic setting makes no such restriction. A transductive learner achieves error $\epsilon=\epsilon(n)$ if its expected loss is at most $\epsilon$ in excess of that incurred by the best fixed predictor in $\mathcal{H}$. As in the PAC setting, the sample complexity $m_{\text{Trans}}(\epsilon)$ is defined inversely.
Similar to the PAC learning, numerous studies have investigated the sample complexity of transductive learning for various domains [20, 19, 21, 28, 4, 3, 7]. Notably, Haussler et al. [20] introduced the one-inclusion-graph (OIG) algorithm for realizable binary classification, which achieves optimal sample complexity (also confirmed in [19, 21]). Subsequently, Rubinstein et al. [28] extended this algorithm to realizable multi-class classification problems, and Bartlett and Long [7] extended the algorithm to partial concept learning. Recently, Asilis et al. [4] extended the OIG approach to agnostic classification, also deriving optimal learners.
Since it demands fine-grained sample-by-sample guarantees, the transductive setting poses additional difficulties beyond those of the distributional setting of PAC learning. On the other hand, transductive learning demands only expected error guarantees, which fall short of the high probability guarantees required for PAC learning. Despite these differences, Warmuth [34] asked whether optimal transductive learners can be used as optimal PAC learners, introducing a new perspective into the study of PAC learning. However, Aden-Ali et al. [2] negatively resolved this question, showing that optimal transductive learners might not ensure high probability guarantees out of the box.
Subsequent work by Aden-Ali et al. [1] revealed that essentially optimal PAC learners for various realizable learning problems could be derived by aggregating the outputs of multiple instances of a transductive learner, trained on different prefixes of the sample. To obtain this result, the authors used a modification of a technique used by Wu et al. [35] to obtain a bound on performance of an online learner. On the other hand, a well-known folklore result (also shown in [4]) demonstrates that realizable PAC learner can be converted into a transductive learner with at most a constant increase in the error rate. These results show that, in realizable supervised learning with a large class of loss functions, PAC learning and transductive learning are essentially equivalent.
The situation is more nuanced in the agnostic setting, which is the focus of the present paper. The naive reduction from transductive to PAC learning suffers blowup of $1/\epsilon$ in the sample complexity (see e.g. [4]), and this is the case even for binary classification. No better reductions are known, to our knowledge. Conversely, prior to our work there were no known general-purpose reductions from PAC to transductive learning in the agnostic setting that guarantee only a constant factor increase in the sample complexity.
1.2 Our Results
This paper tackles the following question in settings where it has remained unanswered.
Are the PAC and transductive models essentially equivalent in a rich variety of supervised learning problems?
We contribute to resolving this question in two ways. Our first result is a framework that reduces PAC learning to transductive learning, building on that of Aden-Ali et al. [1]. As in [1], we train several predictors using the provided transductive learner on prefixes of the same sample, then aggregate their outputs to obtain a PAC learner. Whereas [1] focuses on several prominent loss functions in the realizable setting, our results apply to both realizable and agnostic learning with bounded (pseudo)-metric losses. For our results in the agnostic setting, we need a somewhat more involved analysis using martingales. That said, we only employ off-the-shelf martingale inequalities for both realizable and agnostic analyses. Generalizing to pseudometric losses requires a unified aggregation step. In the realizable setting we use the median in general metric spaces, and the mean for slightly improved constants in metric spaces satisfying Jensen’s inequality. In the agnostic setting, our aggregation step employs a hold-out validation set.
Our second contribution is a characterization of the sample complexity of binary classification in the agnostic transductive setting. This is obtained through analysis of the agnostic OIG algorithm of Asilis et al. [4]. We begin by relating the agnostic sample complexity to the empirical Rademacher complexity of the hypothesis class $\mathcal{H}$ for a given sample $S$. Then, using a well-known upper bound on the Rademacher complexity of finite VC classes [16, 19], we conclude that the agnostic OIG algorithm has sample complexity on the order of $\frac{\operatorname{VC}(\mathcal{H})}{\epsilon^{2}}$, where $\operatorname{VC}(\mathcal{H})$ denotes the VC dimension of the hypothesis class. This result implies that PAC learners obtained from the agnostic OIG algorithm via our reduction achieve essentially optimal PAC sample complexity for the binary classification problem.
In the realizable setting, our first result (modestly) generalizes the approximate equivalence of PAC and transductive learning to problems with pseudometric losses. In the agnostic setting, our reduction shows that PAC learning is essentially no more difficult than transductive learning. However, it remains open whether transductive learning is more difficult than PAC learning due to its fine-grained sample-by-sample requirements. Our second result shows that this is not the case for binary classification, where the agnostic PAC learner derived through our reduction is essentially optimal. We conjecture that this phenomenon might extend to a broader spectrum of agnostic learning problems with pseudometric losses, or even beyond.
2 Technical Overview
2.1 Problem Definition
A supervised learning problem is determined by: a domain $\mathcal{X}$, a label set $\mathcal{Y}$, a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{\geq 0}$, and a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$. For such a task, we provide a finite set of labeled examples $S=\{(x_{1},y_{1}),\dots(x_{n},y_{n})\}\in(\mathcal{X}\times\mathcal{Y})^{n}$ as input to a learning algorithm $A$. This learning algorithm maps the set $S$ to a function from $\mathcal{X}$ to $\mathcal{Y}$, denoted as $A(S)=h$, which we call the learned model or predictor.
[…]
$\ell(y,y^{\prime}):=[y\neq y^{\prime}].$
This definition employs the Iverson bracket notation $[P]$, which equals $1$ if the predicate $P$ is true and 00 otherwise.
In the literature, numerous criteria have been developed to assess the effectiveness of prediction models. Two notable frameworks are the Probably Approximately Correct (PAC) learning model [32], and the transductive learning model [20]. We begin by discussing the PAC learning model introduced by Valiant.
In PAC learning, the error of a predictor $h$, given a probability distribution $\mathcal{D}$ over the sample space $\mathcal{X}\times\mathcal{Y}$, is defined as the expected loss:
$L_{\mathcal{D}}(h)=\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}}[\ell(h(x),y)].$
In addition, we consider a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, against which we benchmark our predictor. Predictors $h\in\mathcal{H}$ are referred to as hypotheses. We will say that $\mathcal{D}$ is realizable by $\mathcal{H}$ if there exists a hypothesis $h^{*}\in\mathcal{H}$ which attains $L_{\mathcal{D}}(h^{*})=0$.
The following definition establishes the concept of PAC learning in the most general terms.
Definition 1 (PAC Learner).
Let $\mathbb{D}$ be a class of probability distributions over $\mathcal{X}\times\mathcal{Y}$ and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. A learner $A$ is a PAC learner for $\mathcal{H}$ with respect to $\mathbb{D}$ if there exists a function $m:(0,1)^{2}\rightarrow\mathbb{Z}_{\geq 1}$ such that for any $\mathcal{D}\in\mathbb{D}$ and $\epsilon,\delta\in(0,1)$, with probability at least $1-\delta$ over the choices of $S$ and any internal randomness in $A$, an i.i.d sample $S$ from $\mathcal{D}$ of size at least $m(\epsilon,\delta)$ satisfies
$L_{\mathcal{D}}(A(S))\leq\mathop{min}_{h\in\mathcal{H}}L_{\mathcal{D}}(h)+\epsilon.$
Moreover, the pointwise minimal sample function $m_{\text{PAC},A}$ of all sample functions for $A$, called the sample complexity of the learner $A$, is defined as
$m_{\text{PAC},A}(\epsilon,\delta)=\mathop{argmin}_{n}\mathop{Pr}_{S\sim\mathcal{D}^{n}}\left[L_{\mathcal{D}}(A(S))\geq\epsilon\right]\leq\delta.$
for all $\mathcal{D}\in\mathbb{D}$.
When $\mathbb{D}$ is the set of all probability distributions realizable by $\mathcal{H}$, we say that $A$ is a realizable PAC learner. We also define the sample complexity of the realizable PAC learning problem for the class $\mathcal{H}$ and the probability distribution $\mathcal{D}$ as the pointwise minimal sample complexity that is attainable by any learner, i.e.,
$m_{\text{PAC},\mathcal{H}}(\epsilon,\delta)=\mathop{min}_{A}m_{\text{PAC},A}(\epsilon,\delta).$
When $\mathbb{D}$ is the set of all probability distributions over $\mathcal{X}\times\mathcal{Y}$, we say that $A$ is an agnostic PAC learner. We define a refined error notion to avoid notation clutter, called agnostic distributional error, as
$L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(A(S)):=L_{\mathcal{D}}(A(S))-\mathop{min}_{h\in\mathcal{H}}L_{\mathcal{D}}(h).$
The criterion for agnostic PAC learnability is met when there exists a learner $A$ such that for any $(\epsilon,\delta)\in(0,1)^{2}$, the agnostic distributional error remains below $\epsilon$ with probability at least $1-\delta$ on samples $S$ with $|S|>n$ for some $n$. Similar to the realizable setup, the sample complexity of the agnostic PAC learning problem for the class $\mathcal{H}$ is the pointwise minimal sample complexity that is needed by any learner, i.e.,
$m^{\text{Ag}}_{\text{PAC},\mathcal{H}}(\epsilon,\delta)=\mathop{min}_{A}m_{\text{PAC},A}^{\text{Ag}}(\epsilon,\delta)$
where, for a learner $A$, $m_{\text{PAC},A}^{\text{Ag}}$ denotes the point-wise minimum of all agnostic sample functions for $A$. For binary classification, the sample complexities are known up to a constant factor. The sample complexity for the realizable case, as established by [17, 23, 1], is given by
$m_{\text{PAC},\mathcal{H}}(\epsilon,\delta)=\Theta\left(\frac{\operatorname{VC}(\mathcal{H})+\log(1/\delta)}{\epsilon}\right),$
while for the agnostic case, as shown by [18], it is
$m_{\text{PAC},\mathcal{H}}^{\text{Ag}}(\epsilon,\delta)=\Theta\left(\frac{\operatorname{VC}(\mathcal{H})+\log(1/\delta)}{\epsilon^{2}}\right).$
In contrast to PAC learning, the transductive learning model evaluates a learner’s performance against an adversarially selected sample $S$. Given samples $S$ of size $n$, let $S_{-i}$ denote the set excluding the $i$-th element, i.e., $S_{-i}=(x_{1},y_{1}),\dots(x_{i-1},y_{i-1}),(x_{i+1},y_{i+1}),\dots(x_{n},y_{n})$. The transductive error, a.k.a. leave-one-out error, for a learner $A$ on a sample $S$ is defined as [20]:
$L^{\text{Trans}}_{S}(A):=\frac{1}{n}\sum_{i\in[n]}\ell(A(S_{-i})(x_{i}),y_{i})$
Analogously, the agnostic transductive error, for a learner $A$ on sample $S$ with respect to a hypothesis class $\mathcal{H}$ is defined as [4]:
$L^{\text{Trans},\text{Ag}}_{S,\mathcal{H}}(A):=L^{\text{Trans}}_{S}(A)-\mathop{min}_{h\in\mathcal{H}}L^{\text{Trans}}_{S}(h).$
We call $S$ realizable if there exists $h^{*}\in\mathcal{H}$ such that $h^{*}_{S}(x_{i})=y_{i}$ for each $(x_{i},y_{i})\in S$. In such a case, the second term disappears, and the agnostic transductive error definition coincides with the traditional definition of Haussler et al. [20].
Definition 2 (Transductive Learner).
Given a sample space $\mathcal{X}\times\mathcal{Y}$ and a hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, a learner $A$ is called a (realizable) transductive learner for $\mathcal{X}\times\mathcal{Y}$ and $\mathcal{H}$, with an error rate function $\epsilon(n)$, if for any sample set $S$ drawn from $(\mathcal{X}\times\mathcal{Y})^{n}$ that is realizable by $\mathcal{H}$,
$L_{S}^{\text{Trans}}(A)\leq\epsilon(n).$
Additionally, $A$ is called an agnostic transductive learner with an error rate function $\epsilon_{\text{Ag}}(n)$, if for any sample set $S$ (not necessarily realizable) drawn from $(\mathcal{X}\times\mathcal{Y})^{n}$,
$L_{S,\mathcal{H}}^{\text{Trans},\text{Ag}}(A)\leq\epsilon_{\text{Ag}}(n).$
Analogous to PAC learning, the sample complexity for a learner $A$ is defined as $m_{\text{Trans},A}(\epsilon^{\prime})=\mathop{min}\{n:\epsilon(m)<\epsilon^{\prime}\ \forall m\geq n\}$. Similarly, one can define the sample complexity of learner $A$ as $m_{\text{Trans},A}^{\text{Ag}}(\epsilon^{\prime})=\mathop{min}\{n:\epsilon_{\text{Ag}}(m)\leq\epsilon^{\prime}\ \forall m\geq n\}$. Furthermore, the transductive sample complexity for the hypothesis class $\mathcal{H}$, for both realizable and agnostic settings, is defined respectively as:
$m_{\text{Trans},\mathcal{H}}(\epsilon)$ $=\mathop{min}_{A}m_{\text{Trans},A}(\epsilon),$
$m^{\text{Ag}}_{\text{Trans},\mathcal{H}}(\epsilon)$ $=\mathop{min}_{A}m^{\text{Ag}}_{\text{Trans},A}(\epsilon).$
Throughout the paper, we make the following two assumptions for transductive learners: (i) a learner $A$ is symmetric with respect to the order of the input, i.e., $A(S)=A(\pi(S))$ for all permutations $\pi$, and (ii) the error rate of any transductive learner, $\epsilon(n)$ or $\epsilon_{\text{Ag}}(n)$, is a non-increasing function of $n$. These assumptions are without loss of generality for randomized learners. For details on why this holds, see Appendix C.
The primary goal of this work is to move toward resolving Warmuth [34]’s conceptual question on constructing optimal PAC learners from transductive learners, moving beyond the scope of realizable binary classification. Our main results include:
  1. 1.
    A framework [^3] Our framework transforms (agnostic) transductive learners to PAC learners for learning problems with bounded pseudometric loss functions. for converting agnostic transductive learners into agnostic PAC learners with comparable error rates, requiring only $\operatorname{poly}(1/\epsilon,\log(1/\delta))$ additional samples, along with a simplification and generalization of the original result of Aden-Ali et al. [1] to arbitrary pseudometric losses.
  2. 2.
    A sharp characterization of the agnostic transductive error rate for binary classification in terms of the VC dimension of the reference hypothesis class. This is obtained through analyzing the agnostic OIG algorithm of [4]. Combined with the first result, this implies that the agnostic OIG algorithm can be converted into a PAC learner for binary classification with essentially optimal sample complexity.
Throughout this paper, when we say a PAC learner or more formally a PAC sample complexity $m_{\text{PAC}}(\epsilon,\delta)$ is essentially optimal, we mean that there exists an absolute constant $c$ and a polynomial $p(\cdot,\cdot)$ such that $m_{\text{PAC}}(\epsilon,\delta)\leq m^{*}_{\text{PAC}}(\epsilon/c,\delta/c)+p(1/\epsilon,1/\delta)$, where $m^{*}_{\text{PAC}}$ is the optimal sample complexity.
2.2 Reduction from PAC learning to Transductive learning
Initially, we revisit an easy consequence of the central result of Aden-Ali et al. [1] for realizable learning, which serves as the starting point for our work.
Theorem 3.
(Equivalent to Aden-Ali et al. [1, Theorem 2.1]) For any realizable learning problem defined by sample space $\mathcal{X}\times\mathcal{Y}$ with a bounded pseudometric loss function $\ell(\cdot,\cdot)\in[0,1]$ and a hypothesis class $\mathcal{H}$, we have
$m_{\text{PAC},\mathcal{H}}(\epsilon,\delta)\leq m_{\text{Trans},\mathcal{H}}(\epsilon/4)+\frac{3}{\epsilon}\cdot\log\left(\frac{2}{\delta}\right).$
In the special cases of classification, partial concept classification, and regression, Theorem 3 is equivalent [^4] The error bound by Aden-Ali et al. [1] and our sample complexity bound are equivalent up to constants, but they rely on different assumptions about the transductive error rates. Aden-Ali et al. [1, Theorem 2.1] assumes $\epsilon(n)\cdot n$ is increasing, while our theorem assumes $\epsilon(n)$ is decreasing. Both assumptions generally hold for randomized realizable transductive learners. Appendix C provides a detailed discussion of these connections. to Aden-Ali et al. [1, Theorem 2.1] up to constants as we show in Appendix C. That said, we prove Theorem 3 from scratch in this paper for three reasons: First, we believe our approach is simpler and uses only off-the-shelf martingale inequalities, in particular the martingale analogue of the multiplicative Chernoff bound outlined in [22]; second, we provide one proof which captures essentially the entirety of supervised learning with (pseudo) metric losses; third, this derivation serves as an accessible warmup for our main result for agnostic learning. Note that Aden-Ali et al. [1, Theorem 2.1] is expressed as a bound on $\epsilon(n)$ parameterized by $m_{\text{PAC},\mathcal{H}}$.
Our main contribution is the following theorem, which, for the first time, demonstrates the reduction of PAC learning to transductive learning within the agnostic learning framework.
Theorem 4.
For any agnostic learning problem defined by sample space $\mathcal{X}\times\mathcal{Y}$ with a bounded pseudometric loss function $\ell(\cdot,\cdot)\in[0,1]$ and a hypothesis class $\mathcal{H}$, we have
$m^{\text{Ag}}_{\text{PAC},\mathcal{H}}(\epsilon,\delta)\leq m^{\text{Ag}}_{\text{Trans},\mathcal{H}}(\epsilon/4)+\frac{8}{\epsilon^{2}}\cdot\log\left(\frac{3}{\epsilon(n)\cdot\delta}\right).$
Our strategy for constructing a PAC learner, applicable to both realizable and agnostic settings, utilizes a transductive learner as a subroutine in a manner similar to [1]. This entails executing the transductive learner on a set of $k\in\operatorname{poly}(1/\epsilon,\log(1/\delta))$ distinct (not necessarily disjoint) samples. Specifically, let $S=\{(x_{1},y_{1}),\dots(x_{n+k},y_{n+k})\}$ be $n+k$ data points, each independently sampled from a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. We define a sequence of sample subsets $S_{i}=\{(x_{1},y_{1}),\dots(x_{n+i},y_{n+i})\}$. Our PAC learners call the transductive learner $A_{\text{Trans}}$ independently with samples $S_{0},S_{1},\dots S_{k-1}$, generating predictors $h_{0},\dots h_{k}$.
Our main technical lemma, analogous to a similar Lemma in [1], states that the average (agnostic) distributional error across predictors $h_{0},\ldots,h_{k}$ is small with high probability. Whereas our proof in the realizable case follows the outline of that of [1], the agnostic case requires an additional (third) martingale argument concerned with the difference between the best-in-class performance and its empirical estimate. This lemma paves the way for an aggregation strategy by carefully combining the outputs from these predictors. In the realizable case and without any additional assumption on the label space or the loss function, we select the generalized median of the predictions — i.e., the prediction minimizing the sum of losses to the other predictions. This generalizes the majority prediction used by [1] for classification to arbitrary metric losses. In the realizable case with a normed vector space of labels, we can instead use the mean as our prediction to improve the error by a constant factor. This generalizes the regression result of [1] to arbitrary metric losses. In the agnostic setup, such aggregation strategies are ineffective, as they can result in a constant blowup in absolute error. Instead, we employ a small validation set of size $O(\frac{1}{\epsilon^{2}}\log\left(\frac{1}{\epsilon\delta}\right)$ to identify the best of our predictors $h_{0},\ldots,h_{k}$.
2.3 Characterizing Transductive sample complexity of agnostic binary classification
In order to characterize the sample complexity of a learner in the PAC setting using our technique, one must first characterize the sample complexity of the underlying learner in the transductive setting. In the case of realizable binary classification, i.e. $\mathcal{Y}=\{+1,-1\}$, Haussler et al. [20] introduced the One Inclusion Graph (OIG) algorithm, a transductive learner for hypothesis class $\mathcal{H}$ with a transductive sample complexity of $O\left(\frac{\operatorname{VC}(\mathcal{H})}{\epsilon}\right)$, where $\operatorname{VC}(\mathcal{H})$ denotes the VC-dimension of $\mathcal{H}$.
The OIG algorithm, takes as input a sample $S$ and constructs a graph $G_{\mathcal{H}_{|S}}=(V,E)$. Here, the vertex set $V=\mathcal{H}_{|S}$ consists of restrictions of $\mathcal{H}$ into $S$, i.e., all possible classifications of samples $S$ by hypothesis class $\mathcal{H}$. The edge set $E$ contains an edge for every pair of restricted hypothesis $h,h^{\prime}$ if they disagree at exactly one data point from $S$. The learner then finds an orientation of $G_{\mathcal{H}_{|S}}$ that minimizes the maximum outdegree of a node, as this procedure also minimizes the loss of the learner on a worst-case sample. Haussler et al. [20] showed that if the edge density is low everywhere simultaneously, orientations with small outdegree can be constructed for any sample $S$ to ensure small error rates. In particular, Haussler et al. [20] showed that when $\mathcal{H}$ has finite VC dimension, then the OIG algorithm attains an optimal error rate of $\epsilon(n)=\frac{\operatorname{VC}(\mathcal{H})}{n}$. Later works derived this error rate using different techniques [20, 19, 21].
Combining this result with Theorem 3, we observe that a PAC learner designed using this reduction attains essentially optimal sample complexity. The following corollary summarizes this discussion.
Corollary 5 (Follows from [1]).
PAC learners, derived from transductive learners, achieve essentially optimal sample complexity in realizable binary classification for a hypothesis class with finite VC dimension.
However, the situation was less clear for the agnostic binary classification problem. Recently, Asilis et al. [4] extended the OIG algorithm to the agnostic setting by constructing a graph $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$ where each vertex represents a possible binary classification of a given sample set $S$, i.e., $V^{\text{Ag}}=\mathcal{Y}^{S}$, and edges exist between classifiers that disagree on a single sample. The resulting graph $G^{\text{Ag}}_{\mathcal{H}_{|S}}$ forms a Boolean hypercube of dimension $|S|=n$.
Similar to the realizable learner, this agnostic learner orients the edges of $G^{\text{Ag}}_{\mathcal{H}_{|S}}$ to minimize a modified outdegree notion that accounts for the representativeness of the hypothesis class. Asilis et al. [4] introduced the “discounted edge density” notion, which effectively determines the best possible error rate achievable by an agnostic transductive learner.
Definition 6.
Given an agnostic one inclusion graph $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$, the discounted edge density of a subset of vertices $U\subseteq V^{\text{Ag}}$ is defined as:
$\Phi_{\text{discounted}}(U):=\frac{|E^{\text{Ag}}(U,U)|-\sum_{u\in U}\|u-\mathcal{H}_{|S}\|_{0}}{|U|},$
where $E^{\text{Ag}}(U,U)$ is the set of edges in $G^{\text{Ag}}_{\mathcal{H}_{|S}}$ for which both endpoints lie in $U$, and $\|u-\mathcal{H}_{|S}\|_{0}$ is the minimum disagreement between the classifier $u$ and a hypothesis from $\mathcal{H}$ on samples $S$.
A familiar reader might recognize that the left term of the ratio in the definition of $\Phi_{\text{discounted}}$ is actually the subgraph edge density introduced by Haussler et al. [20]. In the agnostic setting, the negative term represents how far each particular class assignment is from the optimal hypothesis in the class $\mathcal{H}$. Therefore, we can think of this negative term as “credit” and observe that, all else held equal, larger credit values lead to lower agnostic transductive error rates. In a sense, credit measures how suboptimal this hypothesis class is for general assignments, thereby allowing a learner to compete with $\mathcal{H}$.
The following lemma establishes the optimal error rate of transductive learners for agnostic binary classification, based on the discounted edge density:
Lemma 7.
[Implied by Asilis et al. [4]] The agnostic OIG algorithm attains the optimal agnostic error rate for the transductive learning problem, expressed as
$\epsilon_{\text{Ag}}(n)=\frac{1}{n}\cdot\mathop{max}_{S\in\mathcal{X}^{n}}\mathop{max}_{U\subseteq V^{\text{Ag}}}\Phi_{\text{discounted}}(U)$
where $V^{\text{Ag}}$ denotes the vertex set of $G^{\text{Ag}}_{\mathcal{H}_{|S}}$.
However, this lemma does not provide an explicit characterization of the agnostic transductive error rate in terms of the VC-dimension of $\mathcal{H}$. For the first time, we characterize the performance of the agnostic OIG for agnostic transductive binary classification problem in terms of the VC dimension of the benchmark hypothesis class. In particular, we show the following theorem.
Theorem 8.
The Agnostic One Inclusion Graph algorithm [4] for binary classification with benchmark hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ attains agnostic error rate of $\epsilon(n)=16\cdot\sqrt{\frac{\operatorname{VC}(\mathcal{H})}{n}}$ and sample complexity of at most $O\left(\frac{\operatorname{VC}(\mathcal{H})}{\epsilon^{2}}\right)$.
This theorem together with Theorem 4 leads to a corollary that expands the optimality of the reduction for realizable binary classification, as outlined in Corollary 5, to include the agnostic scenario.
Corollary 9.
PAC learners, derived from transductive learners, achieve essentially optimal sample complexity in agnostic binary classification.
The proof of Theorem 8 consists of two steps and relies on the combinatorial characterization described in Lemma 7. First, we show that for an arbitrary sample set $S$ and its agnostic OIG $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$, the discounted edge density increases when the subset of classifiers $U\subseteq V^{\text{Ag}}$ is grown to a set which is symmetric with respect to the labeling of an arbitrary data point. Recall that the label set is $\mathcal{Y}=\{+1,-1\}$. We define two sets of classifiers $U_{i\rightarrow+}$ and $U_{i\rightarrow-}$ which each contain copies of every classifier $h$ from $U$ with the decision for the $i$-th data point updated to class $+1$ and class $-1$, respectively. The symmetrization of $U$ with respect to the $i$-th data point is defined as $U^{\prime}=U_{i\rightarrow+}\cup U_{i\rightarrow-}$, and we show that the discounted edge density of $U^{\prime}$ is always larger than that of $U$. This observation implies that the discounted edge density is maximized for the complete set of classifiers, $V^{\text{Ag}}$.
In the second part, we estimate the discounted edge density of the complete set of classifiers. By some algebraic manipulations we show that, for any sample $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ of size $n$
$\Phi_{\text{discounted}}(V^{\text{Ag}})=O(n)\cdot\widehat{\Re}_{S}(\mathcal{H}_{|S})$
where ^$\widehat{\Re}_{S}(\mathcal{H}_{|S})$ is the empirical Rademacher of $\mathcal{H}_{|S}$, defined as
^$\widehat{\Re}_{S}(\mathcal{H}_{|S}):=\frac{1}{n}\mathop{\mathbb{E}}_{\sigma\sim\{+1,-1\}^{n}}\left[\sup_{h\in\mathcal{H}_{|S}}\sum_{i=1}^{n}h(x_{i})\cdot\sigma_{i}\right].$
Finally, we utilize the following well-known upper-bound on Rademacher complexity of finite VC dimension hypothesis classes:
Lemma 10.
[Implied by [16] and [19]] For a sample $S\in(\mathcal{X}\times\mathcal{Y})^{n}$ of size $n$ and finite VC dimension hypothesis class $\mathcal{H}$, we have
^$\widehat{\Re}_{S}(\mathcal{H}_{|S})\leq 31\cdot\sqrt{\frac{\operatorname{VC}(\mathcal{H})}{n}}.$
Lemma 10 is a well-known application of the chaining argument introduced by Dudley [16]. Chaining is a technique to establish upper bounds for the expected value of the maximum (or supremum) of a collection of random variables $\{X_{t}\}_{t\in T}$, denoted as $\mathop{\mathbb{E}}[\sup_{t\in T}X_{t}]$. This bound is tighter than the naive sum of their expectations, $\mathop{\mathbb{E}}\left[\sum_{t\in T}X_{t}\right]$, which can be loose when the random variables are highly positively correlated. In such a scenario it makes sense to regroup these random variables that are nearly identical.
When calculating the Rademacher complexity ^$\widehat{\Re}_{S}(\mathcal{H}_{|S})$, the random variables $\left(\sum_{i=1}^{n}h(x_{i})\cdot\sigma_{i}\right)$ are determined by hidden parameters, $h$, forming a metric space under the Hamming distance metric, and their values are close when the hidden parameters are close within the metric space. Dudley’s integral [16] allows us to upper-bound the supremum in terms of the size of the sphere covering of these hidden parameters (hypotheses). On the other hand, Haussler [19] provided an upper bound on the number of spheres that can pack into a hypothesis class $H$ of finite VC dimension that immediately translates to an upper bound on sphere covering of $\mathcal{H}_{|S}$. Combining Dudley’s chaining argument with Haussler’s bound yields this lemma.
For a formal proof of this lemma, see the lecture notes by Rebeschi [27]. For a self-contained introduction to chaining and its applications, we refer readers to Talagrand [31].
3 A simplified reduction in realizable case
In this section, we revisit and extend the technique, originally presented by Aden-Ali et al. [1], for transforming transductive learners to PAC learners for realizable problems. This section serves as a warm-up for our main result: transforming agnostic transductive learners into agnostic PAC learners.
Our method of reduction, in simple terms, diverges from the approach taken by Aden-Ali et al. [1] in terms of problem framing. Their study centers on determining the optimal error bound given a sample of size $n$ by training transductive learners on subsets of the data. In contrast, our strategy explores the requisite additional data needed to transform transductive learners to PAC learners, by training on supersets of the original sample. For a more detailed comparison of the two approaches, see Appendix C.
Data: $n$ i.i.d. samples $S$ from $\mathcal{D}$.
Result: Predictor $\widehat{h}$.
1 Let $A_{\text{Trans}}$ be a transductive learner with error rate $\epsilon(n)$.;
2
3Define $S_{i}:=\{(x_{1},y_{1}),\dots(x_{n},y_{n}),(x_{n+1},y_{n+1}),\dots(x_{n+i},y_{n+i})\}$ for each $i\in[k]$;
4
5Let $h_{i-1}$ be the output predictor of $A_{\text{Trans}}(S_{i-1})$ for each $i\in[k]$.;
6
7Define $\widehat{h}$ as the predictor which returns the median of predictions $h_{0}(x),\dots h_{k-1}(x)$.;
8
9Output: Predictor $\widehat{h}$.;
Algorithm 1 Reduction from PAC Learning to Transductive Learning in the realizable setting.
We streamline the preceding analysis by adopting the following multiplicative variant of Azuma’s inequality rather than formulating a specialized martingale concentration bound.
Lemma 11 (Multiplicative Azuma’s Inequality [22]).
Let $X_{1},\dots X_{n}\in[0,c]$ be real-valued random variables for some $c>0$. Suppose that $\mathop{\mathbb{E}}[X_{i}\mid X_{1},\dots X_{i-1}]\leq a_{i}$ for all $i$. Let $\mu=\sum_{i=1}^{n}a_{i}$. Then for any $\delta>0$,
$\mathop{\mathbf{Pr}}\left[\sum_{i}X_{i}\geq(1+\delta)\cdot\mu\right]\leq\exp\left(-\frac{\delta^{2}\cdot\mu}{(2+\delta)\cdot c}\right).$
Beyond simplifying the process, we have expanded the scope of our reduction to cover any supervised learning problem with bounded metric loss. This class of problems includes a wide array of supervised learning tasks, such as classification, partial concept classification, and bounded regression. For this extension, we have adapted the aggregation phase of the reduction to output the median of the prediction in the metric space, defining the median as the point that minimizes the overall distance to all other predictions.
The following lemma is the key assertion that will allow us to finalize the proof of Theorem 3.
Lemma 12.
Let $A_{\text{Trans}}$ be a transductive learner with transductive error rate $\epsilon(n)$. Then, for any $\delta\in(0,1)$, and a sample $S\sim\mathcal{D}^{n+k}$, with probability $1-\delta$ over the choices of $S$, predictors $h_{i-1}:=A_{\text{Trans}}(S_{i-1})$ for $i\in[k]$ satisfy:
$\frac{1}{k}\sum_{i\in[k]}L_{\mathcal{D}}(h_{i-1})\leq 4\cdot\epsilon(n)$
where $k=\frac{3\cdot\log(2/\delta)}{\epsilon(n)}$.
Before we prove the lemma, let us show that how it implies Theorem 3.
Proof of Theorem 3.
Fix any test point $(x,y)$, let $\widehat{h}$ be the predictor that outputs the median among predictions $h_{0}(x),\dots,h_{k-1}(x)$ which is the label $\widehat{y}$ that minimizes the total loss between $\widehat{y}$ and predictions $\{h_{i-1}(x)\}_{i\in[k]}$.
Thus,
[…]
$\leq\frac{1}{k}\sum_{i\in[k]}\ell(h_{i-1}(x),\widehat{y})+\ell(h_{i-1}(x),y)$ $(\mathop{min}\leq\mathop{avg})$
$\leq 2\cdot\frac{1}{k}\sum_{i\in[k]}\ell(h_{i-1}(x),y).$ $(\widehat{y}\text{ is median})$
Finally, taking expectation over samples $(x,y)\sim\mathcal{D}$ together with Lemma 12 concludes that distributional error of predictor $\widehat{h}$ is at most $8\cdot\epsilon(n)$ with probability $1-\delta$.
When $(\mathcal{Y},\ell)$ constitutes a normed vector space, the application of Jensen’s inequality ensures that the average of the predictors’ outputs, $h_{0}(x),\dots$ $h_{k-1}(x)$ yields at most $4\cdot\epsilon(n)$ distributional error, thereby improves general metric space result. ∎
We now proceed to prove Lemma 12. To begin, let us define the empirical error of predictor $h_{i-1}$ for any $i\in[k]$ against the data point $(x_{n+i},y_{n+i})$ as
$d_{i}:=\ell(h_{i-1}(x_{n+i}),y_{n+i}).$
Given the sample set $S_{i-1}$, $d_{i}$ serves as an unbiased estimate of the distributional error associated with predictor $h_{i-1}$. Formally,
$\mathop{\mathbb{E}}[d_{i}\mid S_{i-1}]=\mathop{\mathbb{E}}[\ell(h_{i-1}(x_{n+i}),y_{n+i})\mid S_{i-1}]=L_{\mathcal{D}}(h_{i-1})$
as we know that $(x_{n+i},y_{n+i})$ is independently sampled from $\mathcal{D}$.
Next, we show that, with high probability, the total distributional error is at most a constant factor larger than the total empirical error. The key observation is that the cumulative differences of $\mathcal{L}_{\mathcal{D}}(h_{i-1})-d_{i}$ form a martingale, and the sequence can be bounded using Azuma’s multiplicative inequality. This concept is detailed in the subsequent claim, whose proof is postponed to Appendix A.1:
Claim 13 (Forward Martingale Bound).
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}L_{\mathcal{D}}(h_{i-1})-\sum_{i=1}^{k}d_{i}>2\cdot k\cdot\epsilon(n)\right]\leq\frac{\delta}{2}.$
Following that, we turn our attention to show that total empirical error is small with high probability. We observe that $d_{i}$ acts as an unbiased estimator for the transductive error of the learner $A_{\text{Trans}}$ given the unordered set $S_{i}$ as input. We then think of constructing our sequence of samples $S_{0},\dots,S_{k}$ “backwards” by first conditioning on the (unordered) set $S_{k}$, then iteratively peeling off one sample at a time — uniformly at random without replacement — to obtain $S_{k-1},S_{k-2},\ldots,S_{0}$. For $i=k,\ldots,1$, observe that $(x_{i},y_{i})$ is a uniformly random element from $S_{i}$, and that $|S_{i}|\geq n$. The transductive error assumption implies that $L_{S_{i}}^{\text{Trans}}(A)\leq\epsilon(n)$. Finally, applying the multiplicative version of Azuma’s inequality a second time, we obtain the following bound. The proof of this bound is deferred to Appendix A.1.
Claim 14 (Backward Martingale Bound).
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}d_{i}>2\cdot k\cdot\epsilon(n)\right]\leq\frac{\delta}{2}.$
Combining two claims by using union bound, we conclude the proof of Lemma 12, from which Theorem 3 follows as previously described.
4 Proof of Theorem 4: Reduction for Agnostic Learning
This section is devoted to the main result. Here, we will prove Theorem 4 by following a proof path similar to realizable setup. In the agnostic case, our goal is to show that the average performance of a trained predictor is not significantly worse than that of the best hypothesis in $\mathcal{H}$. To achieve this goal, we employ forward and backward martingale bounds, akin to those used in the realizable case, to show that the average distributional agnostic loss is not substantially worse than the average observed agnostic loss with high probability. However, the agnostic case necessitates an additional backward martingale bound to establish that sample optimal hypotheses do not perform considerably worse than the distributional optimal hypothesis on their respective sample points.
If the objective is to obtain a randomized predictor that achieves a small distributional loss with high probability over the data distribution but in expectation over its internal randomness, then we can simply select a hypothesis at random and return its prediction on each new data point encountered. Alternatively, if the goal is to obtain a deterministic learner that achieves small distributional loss with high probability over the data distribution, then we utilize a cross-validation set containing a small number of additional data points. By testing each predictor on this cross validation set, we can assert with high probability that the deterministic predictor which performs the best on this cross validation set will also attain a small agnostic distributional loss.
Lemma 15.
Let $A_{\text{Trans}}$ be a transductive learner with agnostic error rate $\epsilon_{\text{Ag}}(n).$ Then for any $\delta\in(0,1)$, and a sample $S\sim\mathcal{D}^{n+k}$, with probability $1-\delta$ over the choices of $S$, predictors $h_{i-1}:=A_{\text{Trans}}(S_{i-1})$ for $i\in[k]$ satisfy:
$\mathop{\mathbf{Pr}}\left[\frac{1}{k}\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})\geq\epsilon_{\text{Ag}}(n)+\sqrt{\frac{72\cdot\log(3/\delta)}{k}}\right]\leq\delta.$
In particular, when $k=\left\lceil{\frac{72\cdot\log(3/\delta)}{\epsilon_{\text{Ag}}^{2}(n)}}\right\rceil$, average distributional error is no more than $2\cdot\epsilon_{\text{Ag}}(n)$ with probability at most $1-\delta$.
In order to show that PAC learning reduces to transductive learning in the agnostic setting, we will use the following algorithm to convert an arbitrary transductive learner to a PAC learner.
Data: $n$ i.i.d. samples $S$ from $\mathcal{D}$.
Result: Predictor $\widehat{h}$.
1 Let $A_{\text{Trans}}$ be a transductive learner with agnostic error rate $\epsilon_{\text{Ag}}(n)$.
2Define $S_{i}:=(x_{1},y_{1}),\dots(x_{n},y_{n}),(x_{n+1},y_{n+1}),\dots(x_{n+i},y_{n+i})$ for each $i\in[k]$.
3Let $h_{i-1}$ be the output predictor of $A_{\text{Trans}}(S_{i-1})$ for each $i\in[k]$.
4Let $S_{\text{Val}}$ be a hold-out validation set of size $k^{\prime}=O\left(\frac{\log(k/\delta)}{\epsilon_{\text{Ag}}^{2}(n)}\right)$.
5Estimate the empirical error of each predictor $\{h_{i-1}\}_{i\in[k]}$ by using validation set $S_{\text{Val}}$
6Set $\widehat{h}$ as the predictor with minimum validation error among $\{h_{i-1}\}_{i\in[k]}$.
Output: Predictor $\widehat{h}$.
Algorithm 2 Transforming transductive learners to PAC learners in the agnostic setting
Before proving Lemma 15, we argue briefly how it implies Theorem 4. Lemma 15 guarantees that the average, and hence also the minimum, of the agnostic risks of $\{h_{0},\ldots,h_{k-1}\}$ is bounded by $2\cdot\epsilon(n)$ with probability $1-\frac{\delta}{2}$. We then use a hold-out validation dataset to select some $h_{i}$ whose agnostic risk is within $\epsilon(n)$ of this bound with probability $1-\frac{\delta}{2}$. Standard Hoeffding bounds and the union bound imply that a validation set of size $k^{\prime}=O(\frac{1}{\epsilon^{2}(n)}\log(\frac{k}{\delta}))$ suffices. Therefore, in total $k+k^{\prime}=O\left(\frac{1}{\epsilon^{2}(n)}\log(\frac{1}{\epsilon(n)\cdot\delta})\right)$ samples suffice, as claimed in Theorem 4. We omit the details of the argument here, as it follows a standard Hoeffding bound argument. Please see Appendix B for the proof details.
We now proceed to prove Lemma 15. Our proof follows a similar trajectory to realizable setup. Let $h^{*}_{\mathcal{D}}$ be the optimal hypothesis in $\mathcal{H}$ for distribution $\mathcal{D}$, i.e., $h^{*}_{\mathcal{D}}\in\mathop{argmin}_{h\in\mathcal{H}}L_{\mathcal{D}}(h)$. Next, we define the following random variable similar to realizable setting:
$d_{i}=\ell(h_{i-1}(x_{i}),y_{i})-\ell(h^{*}_{\mathcal{D}}(x_{i}),y_{i}).$
Also, conditioned on samples $S_{i-1}$, $d_{i}$ becomes an unbiased estimator of the distributional agnostic risk of the predictor $h_{i-1}$. Formally,
$\mathop{\mathbb{E}}[d_{i}\mid S_{i-1}]$ $=\mathop{\mathbb{E}}[\ell(h_{i-1}(x_{i})\neq y_{i})\mid S_{i-1}]-\mathop{\mathbb{E}}[\ell(h_{\mathcal{D}}^{*}(x_{i})\neq y_{i})\mid S_{i-1}]$
$=L_{\mathcal{D}}(h_{i-1})-L_{\mathcal{D}}(h^{*}_{\mathcal{D}})$
$=L_{\mathcal{D},\mathcal{H}}^{ag}(h_{i-1}).$
The proof of Lemma 15 employs both forward and backward martingale bounds, mirroring the approach used in the realizable context. In the forward martingale phase, we aim to demonstrate that $\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})\lesssim\sum_{i=1}^{k}d_{i}$, with high probability, by constructing a martingale and using Azuma’s martingale concentration inequality to show that martingale stays small with high probability.
The backward martingale phase, however, diverges from the realizable scenario. Unlike in the realizable case where $d_{i}$ serves as an unbiased estimator of the transductive error under certain conditions, such a straightforward relationship does not exist here. To navigate this complexity, we introduce $\widetilde{d}_{i}$ as an unbiased estimator for the agnostic transductive error of $A_{\text{Trans}}$ using the sample sets $S_{i}={(x_{1},y_{1}),\dots,(x_{n+i},y_{n+i})}$. It is straightforward to confirm that $\mathop{\mathbb{E}}[d_{i}]\leq\mathop{\mathbb{E}}[\widetilde{d}_{i}]$. However, to establish a probabilistic guarantee, we construct two additional martingales, showing that the series $\sum_{i}\widetilde{d}_{i}$ and the differences $\sum_{i}d_{i}-\widetilde{d}_{i}$ both act as super-martingales. Using Azuma’s inequality twice, we argue that $\sum_{i}d_{i}$ remains small with high probability, thereby completing our proof.
In the agnostic case, the multiplicative version of Azuma’s inequality which we used in the realizable case is not as powerful. Applying it in the agnostic setting could lead to a high loss for our hypothesis class $\mathcal{H}$, and any multiplicative blowup on the martingales defined below would also amplify the error of the reference hypothesis class. This amplification would undermine the possibility of a general PAC learnability guarantee. To that end, in the following concentration bounds, we employ the traditional additive version of Azuma’s inequality in order to obtain the desired guarantees on the error of our predictors.
Claim 16 (Forward Martingale Bound).
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})-\sum_{i=1}^{k}d_{i}>\sqrt{8k\log\left(\frac{3}{\delta}\right)}\right]\leq\delta/3.$
Proof.
Let $M_{i}=L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})-d_{i}$, and $\overline{M}_{i}=\sum_{j=1}^{i}M_{j}$. Then, consider the following conditional probability to see that $\overline{M}_{i}$ forms a martingale.
$\mathop{\mathbb{E}}_{(x_{i},y_{i})\sim\mathcal{D}}[\overline{M}_{i}\mid S_{i-1}]$ $=\mathop{\mathbb{E}}[M_{i}\mid S_{i-1}]+\mathop{\mathbb{E}}[\overline{M}_{i-1}\mid S_{i-1}]$
$=\mathop{\mathbb{E}}[L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})-d_{i}\mid S_{i-1}]+\overline{M}_{i-1}$
$=\overline{M}_{i-1}.$
Since both $d_{i}$ and $L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})$ are bounded between $[-1,1]$, $M_{i}$ lies in $[-2,2]$ and so $|M_{i}|\leq 2$. Applying Azuma’s inequality yields the desired result. ∎
Now, we know that $\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})$ is not much larger than $\sum_{i=1}^{k}d_{i}$, i.e.,
$\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i-1})\lesssim\sum_{i=1}^{k}d_{i}$
with high probability. Next, we show that right hand side is small with high probability by constructing two backward martingales.
For a given sample $S$, let $h_{S}^{*}$ be an optimal hypothesis for this sample. We define the following unbiased estimator of the $i$-th agnostic transductive error $L_{S_{i},\mathcal{H}}^{\text{Trans},\text{Ag}}(A)$
$\widetilde{d}_{i}=\ell(h_{i-1}(x_{i}),y_{i})-\ell(h^{*}_{S_{i}}(x_{i})\neq y_{i}).$
We then decompose the total distributional agnostic error of our predictors as a the sum of transductive errors and the difference between distributional and transductive errors, as follows.
$\sum_{i=1}^{k}d_{i}=\sum_{i=1}^{k}\widetilde{d}_{i}+\sum_{i=1}^{k}(d_{i}-\widetilde{d}_{i})$
In the next two claims, we show that $\sum_{i=1}^{k}d_{i}$ is small with high probability by showing that both component sums are small with high probability. We start by showing that the sum of $\widetilde{d}_{i}$ is small with high probability.
Claim 17 (Backward Martingale Bound I: Total Transductive Error).
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}\widetilde{d}_{i}>k\cdot\epsilon+\sqrt{8k\log\left(\frac{3}{\delta}\right)}\right]\leq\delta/3$
Proof.
We think of constructing our sequence of samples $S_{0},\dots,S_{k}$ “backwards” by first conditioning on the (unordered) set $S_{k}=\{(x_{i},y_{i})\}_{i=1}^{n+k}$, then iteratively peeling off one sample at a time — uniformly at random without replacement — to obtain $S_{k-1},S_{k-2},\ldots,S_{0}=S$. For $i=k,\ldots,1$, observe that $(x_{i},y_{i})$ is a uniformly random element from $S_{i}$, and that $|S_{i}|\geq n$. The transductive error assumption implies that $L_{S_{i},\mathcal{H}}^{\text{Trans},\text{Ag}}(A)\leq\epsilon$ and therefore
$\mathop{\mathbb{E}}[\widetilde{d}_{i}\mid S_{i}]\leq\epsilon.$
Since $\widetilde{d}_{i}$ is conditionally independent of $\widetilde{d}_{k},\ldots,\widetilde{d}_{i+1}$ given $S_{i}$, it follows that
$\mathop{\mathbb{E}}[\widetilde{d}_{i}\mid\widetilde{d}_{k},\ldots,\widetilde{d}_{i+1}]\leq\epsilon.$
Next, we define random variables $B_{i}=\widetilde{d}_{i}-\epsilon$, $\overline{B}_{i}=\sum_{j=i}^{k}B_{i}$ and observe that
$\mathop{\mathbb{E}}[\overline{B_{i}}\mid\overline{B}_{i+1}]=\overline{B}_{i+1}+\mathop{\mathbb{E}}[\widetilde{d}_{i}-\epsilon\mid\widetilde{d}_{k},\ldots,\widetilde{d}_{i+1}]\leq\overline{B}_{i+1}$
Thus, the sequence $\overline{B}_{i}$ is a backwards super-martingale — i.e., a super-martingale when viewed backwards in time. Invoking Azuma’s inequality, we obtain
$\mathop{\mathbf{Pr}}\left[\overline{B}_{1}>k\cdot\epsilon(n)\right]=\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}\widetilde{d}_{i}>2\cdot k\cdot\epsilon(n)\right]\leq\frac{\delta}{3}.$
Next, we investigate the second sum. By definition, $h^{*}_{S_{i}}$ is the optimal hypothesis for $S_{i}$, which implies that its expected performance on that sample is at least as good as that of $h^{*}_{\mathcal{D}}$. In this bound, we strengthen the in-expectation guarantee to a high probability guarantee, showing that the sequence of $h^{*}_{S_{i}}$ predictors will not perform significantly worse on their respective test points compared to $h^{*}_{\mathcal{D}}$.
Claim 18 (Backward Martingale Bound II: Transductive vs Distributional Error).
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}(d_{i}-\widetilde{d}_{i})>\sqrt{8k\log\left(\frac{3}{\delta}\right)}\right]<\frac{\delta}{3}.$
We decided to leave the proof of this claim for the appendix, as it follows a similar structure to the previous proofs.
Combining Claims 16, 17, and 18 using the union bound, we obtain
$\mathop{\mathbf{Pr}}\left[\sum_{i=1}^{k}L_{\mathcal{D},\mathcal{H}}^{ag}(h_{i-1})>k\cdot\epsilon+3\sqrt{8k\log\left(\frac{3}{\delta}\right)}\right]\leq\delta$
and therefore,
$\mathop{\mathbf{Pr}}\left[\frac{1}{k}\sum_{i=0}^{k-1}L_{\mathcal{D},\mathcal{H}}^{ag}(h_{i})>\epsilon+\sqrt{\frac{72\cdot\log(3/\delta)}{k}}\right]\leq\delta.$
This completes the proof of Lemma 15, from which Theorem 4 follows as previously described.
5 Optimal Agnostic Transductive Learner for Binary Classification
In this section, we aim to establish an agnostic transductive error rate of binary classification in terms of the VC dimension of benchmark hypothesis class $\mathcal{H}$. We analyze the agnostic OIG algorithm of Asilis et al. [4] and show that together with our reduction, the agnostic OIG algorithm obtains essentially optimal PAC sample complexity. The main result of this section is restated here for the clarity of the discussion. See 8
Before delving into the proof of this theorem, it is essential to discuss its implications. Theorem 8 suggests that the sample complexity of agnostic transductive learning for binary classification is at most $O\left(\frac{\operatorname{VC}(\mathcal{H})}{\epsilon^{2}}\right)$. Let $A_{\text{Trans}}$ denote the agnostic transductive learner guaranteed by this theorem. By invoking Theorem 4, we can infer that Algorithm 2 transforms $A_{\text{Trans}}$ into a PAC learner with a sample complexity of $O\left(\frac{\operatorname{VC}(\mathcal{H})+\log(1/(\epsilon\delta))}{\epsilon^{2}}\right)$. Conversely, Theorem 4 and the fundamental theorem of statistical learning theory [29] collectively imply a lower bound of $\Omega\left(\frac{\operatorname{VC}(\mathcal{H})+\log(1/\delta)}{\epsilon^{2}}\right)$ for the error rate of an agnostic transductive learner. Consequently, when $\epsilon=O(\operatorname{poly}(\delta))$, which is generally true in practice, we see that our theorem is optimal up to a constant factor.
To commence our analysis, we introduce the agnostic one inclusion graph algorithm, as proposed by Asilis et al. [4]. Consider an agnostic binary classification problem with a benchmark hypothesis class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, where $\mathcal{Y}=\{+1,-1\}$, and let $S\subseteq(\mathcal{X}\times\mathcal{Y})^{n}$ be a sample of size $n$. The agnostic one inclusion graph (OIG) $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$ is defined by the following vertex and edge sets:
  • $V^{\text{Ag}}=\mathcal{Y}^{n}$, representing one node for each possible labeling of the $n$ data points.
  • $E^{\text{Ag}}=\{(u,v):\|u-v\|_{0}=1\}$, where the Hamming distance $\|u-v\|_{0}$ is defined as $\sum_{i=1}^{n}[u_{i}\neq v_{i}]$, representing the number of indices at which $u$ and $v$ differ.
It is worth noting that $G^{\text{Ag}}_{\mathcal{H}_{|S}}$ is isomorphic to the Boolean hypercube of dimension $|S|=n$.
The agnostic one inclusion graph algorithm, similarly to its realizable counterpart, orients the edges of the agnostic OIG to ensure that each vertex has a small out-degree. The labeling of $n-1$ data points (training data) corresponds to an edge of this agnostic OIG, and the algorithm selects the vertex along the direction of this edge, outputting the label of the test data according to this vertex. Now, we recall the “discounted edge density” notion of agnostic OIGs and restate a lemma from [4] which identifies optimal error rate of agnostic OIG algorithm based on maximum local discounted edge density.
$\Phi_{\text{discounted}}(U):=\frac{|E(U,U)|-\sum_{u\in U}\|u-\mathcal{H}_{|S}\|_{0}}{|U|}$
where $\|v-U\|_{0}$ defines the Hamming distance between a vertex $v$ and a subset $U$ as the minimum Hamming distance between $v$ and the members of $U$, i.e., $\|v-U\|_{0}=\mathop{min}_{u\in U}\|v-u\|_{0}$. The following lemma characterizes the optimal agnostic transductive sample complexity of the agnostic OIG algorithm and therefore, the binary classification problem as shown by Asilis et al. [4].
See 7
In the rest of this section, we utilize Asilis et al. [4]’s characterization to prove Theorem 8, which states that the agnostic OIG algorithm admits an $\epsilon(n)=16\cdot\sqrt{\frac{\operatorname{VC}(\mathcal{H})}{n}}$ error rate. Our proof consists of two parts. First, we show that the expression given in the lemma is maximized when $U=V^{\text{Ag}}$, or equivalently, for the complete Boolean hypercube. The following lemma formalizes this argument.
Lemma 19.
Given a sample $S$ and hypothesis class $\mathcal{H}$, let $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$ be their agnostic OIG. The following inequality holds for all $U\subseteq V^{\text{Ag}}$:
$\Phi_{\text{discounted}}(U)\leq\Phi_{\text{discounted}}(V^{\text{Ag}}).$
The second part of the proof is to bound this ratio for a maximal instance where $U=V^{\text{Ag}}$. We consider the hypothesis class $\mathcal{H}$ and the sample set $S$ to be fixed arbitrarily, and aim to prove the following lemma.
Lemma 20.
Given a sample $S$ and hypothesis class $\mathcal{H}$, let $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$ be their agnostic OIG. Then, we have $\Phi_{\text{discounted}}(V^{\text{Ag}})\leq 16\cdot\sqrt{n\cdot\operatorname{VC}(\mathcal{H})}.$
Before presenting the proofs of these lemmas, we first prove the main result of this section by using them. For any arbitrary sample set $S\subseteq(\mathcal{X}\times\mathcal{Y})^{n}$ of size $n$, let $G^{\text{Ag}}_{\mathcal{H}_{|S}}=(V^{\text{Ag}},E^{\text{Ag}})$ be the agnostic OIG for $S$, Lemma 19 and Lemma 20 imply that
$\mathop{max}_{U\subseteq V}\Phi_{\text{discounted}}(U)=\Phi_{\text{discounted}}(V^{\text{Ag}})\leq 16\cdot\sqrt{n\cdot\operatorname{VC}(\mathcal{H})}.$
Since $S$ is an arbitrary sample set, invoking Lemma 7 ensures that the error rate is at most ${16\cdot\sqrt{\frac{\operatorname{VC}(\mathcal{H})}{n}}}$, completing the proof of Theorem 8. The remaining part of this section is devoted to the proofs of Lemma 19 and Lemma 20.
5.1 Proof of Lemma 19
To avoid notational clutter, we omit the superscript “Ag” for $V^{\text{Ag}}$ and $E^{\text{Ag}}$ throughout the proof, denoting them simply as $V$ and $E$, respectively. We first define the total credit (or total Hamming distance) for any subset $U\subseteq V$ as $\|U-\mathcal{H}_{|S}\|_{0}:=\sum_{u\in U}\|u-\mathcal{H}_{|S}\|_{0}$. We fix an arbitrary sample size $|S|=n$ for some $n\geq 3$. Note that each vertex $v\in V$ denotes a label assignment for samples in $S$. We define $v_{i\rightarrow+}$ as another vertex such that $i^{\text{th}}$ sample is assigned to class label $+1$.
Formally,
$v_{i\rightarrow+}(x_{j})=\begin{cases}v(x_{j})&\text{for any }(x_{j},y_{j})\in S_{-i}\\ +1&j=i.\end{cases}\qquad v_{i\rightarrow-}(x_{j})=\begin{cases}v(x_{j})&\text{for any }(x_{j},y_{j})\in S_{-i}\\ -1&j=i.\end{cases}$
We extend this definition to subsets of vertices as $U_{i\rightarrow+}:=\{u_{i\rightarrow+}:u\in U\}$ and $U_{i\rightarrow-}:=\{u_{i\rightarrow-}:u\in U\}$.
Our objective is to demonstrate that $\Phi_{\text{discounted}}(V)\geq\Phi_{\text{discounted}}(U)$ for any $U\subseteq V$. Let $\alpha$ be the maximum possible value of the function $\Phi_{\text{discounted}}$ attained at any subset, i.e., $\alpha=\mathop{max}_{U\subseteq V}\Phi_{\text{discounted}}(U)$, and let $U$ be an arbitrary non-empty set satisfying $\Phi_{\text{discounted}}(U)=\alpha$. As $\Phi_{\text{discounted}}$ is positive for a subset $U$ that contains two vertices: a hypothesis $h\in\mathcal{H}_{|S}$ and its arbitrary neighbor, $\alpha>0$ always and $U$ is non-empty. Next, we define
$g(T):=|E(T,T)|-\|T-\mathcal{H}_{|S}\|_{0}-\alpha\cdot|T|.$
Recall that $E(T,T)$ is the collection of edges with both endpoints in $T$. Our goal is to show that $g(V)\geq 0$. Notice that $g(V)\geq 0$ implies $\Phi_{\text{discounted}}(V)\geq\alpha$.
Fix an arbitrary index $i\in[n]$ and data point $(x_{i},y_{i})\in S$. Let $V^{+}$ and $V^{-}$ form a partition of $V$ such that $V^{+}:=\{v\in V:v(x_{i})=+1\}$ and $V^{-}:=\{v\in V:v(x_{i})=-1\}$. By using the sets $V^{+}$ and $V^{-}$, we also define a partition of $U$ as $U^{+}=U\cap V^{+}$ and $U^{-}=U\cap V^{-}$. We further define the following sets:
$N^{+}=U^{+}\cap U^{-}_{i\rightarrow+}$ $N^{-}=U^{-}\cap U^{+}_{i\rightarrow-}$
$I^{+}=U^{+}\setminus N^{+}$ $I^{-}=U^{-}\setminus N^{-}.$
Here, $N^{+},I^{+},N^{-},I^{-}$ form a partition of $U$. The intuition behind this partitioning is to analyze the behavior of the function $g$ when the $i$-th bit is flipped for each part of the set $U$. We refer the reader to Figure 1 for a visualization of these sets.
By examining the change in the value of $g$ under this bit-flipping operation, we aim to establish a relationship between $g(U)$ and $g(U^{\prime})$ where $U^{\prime}$ is obtained by adding the minimal set of vertices to $U$ which yields a symmetric set $U^{\prime}$ with respect to $i$-th coordinate, i.e. $U^{\prime}=U_{i\rightarrow 0}\cup U_{i\rightarrow 1}$. Ultimately, this will lead to the desired inequality $\Phi_{\text{discounted}}(V)\geq\Phi_{\text{discounted}}(U)$. Next, we observe that
$g(U)=g(I^{+})+g(I^{-})+g(N^{+}\cup N^{-})+|E(I^{+},N^{+})|+|E(I^{-},N^{-})|.$
[…]
Figure 1: Symmetrization argument: A Boolean hypercube of dimension $4$ is partitioned into two smaller hypercubes of dimension $3$ by fixing a certain coordinate $i$. The smaller hypercubes $V^{+}$ and $V^{-}$ are visualized as $8$ boxes in the figure. Vertically aligned boxes correspond to bit strings that differ only at index $i$. The set $U$ is indicated by the blue boxes.
Let us define another set $U^{\prime}:=U\cup I^{+}_{i\rightarrow-}\cup I^{-}_{i\rightarrow+}$ and evaluate $g$’s value at $U^{\prime}$ as
$g(U^{\prime})$ $=g(U)+g(I^{+}_{i\rightarrow-})+g(I^{-}_{i\rightarrow+})$
$+|E(I^{+}_{i\rightarrow-},U^{-})|+|E(I^{-}_{i\rightarrow+},U^{+})|$ (1)
$+|E(I^{+}_{i\rightarrow-},I^{+})|+|E(I^{-}_{i\rightarrow+},I^{-})|.$
Now, we make series of observations:
  1. 1.
    For each vertex $v\in V$, because $\|v_{i\rightarrow-}-v_{i\rightarrow+}\|_{0}=1$ we see by the triangle inequality that
    $\mathop{max}\{\|v_{i\rightarrow+}-\mathcal{H}_{|S}\|_{0},\|v_{i\rightarrow-}-\mathcal{H}_{|S}\|_{0}\}\leq\mathop{min}\{\|v_{i\rightarrow+}-\mathcal{H}_{|S}\|_{0},\|v_{i\rightarrow-}-\mathcal{H}_{|S}\|_{0}\}+1.$
    Therefore, we have
    $\|v_{i\rightarrow+}-\mathcal{H}_{|S}\|_{0}+\|v_{i\rightarrow-}-\mathcal{H}_{|S}\|_{0}\leq 2\cdot\mathop{min}\{\|v_{i\rightarrow+}-\mathcal{H}_{|S}\|_{0},\|v_{i\rightarrow-}-\mathcal{H}_{|S}\|_{0}\}+1.$
  2. 2.
    From the first observation we deduce that
    $\|I^{+}_{i\rightarrow-}-\mathcal{H}_{|S}\|_{0}\leq\|I^{+}-\mathcal{H}_{|S}\|_{0}+|I^{+}|\qquad\text{and}\qquad\|I^{-}_{i\rightarrow+}-\mathcal{H}_{|S}\|_{0}\leq\|I^{-}-\mathcal{H}_{|S}\|_{0}+|I^{-}|.$
    Since $|E(I^{+}_{i\rightarrow-},I^{+}_{i\rightarrow-})|=|E(I^{+},I^{+})|$ and $|E(I^{-}_{i\rightarrow+},I^{-}_{i\rightarrow+})|=|E(I^{-},I^{-})|$, subtracting these quantities from the previous inequalities, we see that
    $g(I^{+}_{i\rightarrow-})\geq g(I^{+})-|I^{+}|\qquad\text{and}\qquad g(I^{-}_{i\rightarrow+})\geq g(I^{-})-|I^{-}|$
  3. 3.
    For edge sets we observe
    $|E(I^{+}_{i\rightarrow-},U^{-})|\geq|E(I^{+},N^{+})|\qquad\text{and}\qquad|E(I^{-}_{i\rightarrow+},U^{+})|\geq|E(I^{-},N^{-})|$
    and also
    $|E(I^{+}_{i\rightarrow-},I^{+})|=|I^{+}|\qquad\text{and}\qquad|E(I^{-}_{i\rightarrow+},I^{-})|=|I^{-}|.$
By substituting these inequalities to (5.1) we obtain that
$g(U^{\prime})-g(U)$ $\geq g(I^{+})-|I^{+}|+g(I^{-})-|I^{-}|+|E(I^{+},N^{+})|+|E(I^{-},N^{-})|+|I^{+}|+|I^{-}|$
$=g(I^{+})+g(I^{-})+|E(I^{+},N^{+})|+|E(I^{-},N^{-})|$
$=g(U)-g(N^{+}\cup N^{-}).$
We claim that $g(U)-g(N^{+}\cup N^{-})$ is non-negative since assuming otherwise implies that $g(N^{+}\cup N^{-})>0$ and so $\Phi_{\text{discounted}}(N^{+}\cup N^{-})>\alpha$ which contradicts with the optimality of $U$. Thus, $g(U^{\prime})\geq g(U)$, and so $\Phi_{\text{discounted}}(U^{\prime})\geq\Phi_{\text{discounted}}(U)$.
Note that by construction of the set $U^{\prime}$, it is the minimal superset of $U$ which is symmetric with respect to the index $i$. Therefore, by repeating this process for each index $i\in[n]$, to create a set $U^{\prime\prime}$, $g$ will only increase. This implies that the maximal subset $U^{\prime\prime}$ that satisfies $g(U^{\prime\prime})\geq 0$ or $\Phi_{\text{discounted}}(U^{\prime\prime})\geq\alpha$ is symmetric with respect to any index. The only non-empty subset satisfying this property is $V$ itself, the complete boolean hypercube. Therefore, the proof is complete.
5.2 Proof of Lemma 20
We first restate the lemma here: See 20
Proof.
To simplify notation, we omit the superscript “Ag” for $V^{\text{Ag}}$ and $E^{\text{Ag}}$ throughout the proof, denoting them as $V$ and $E$, respectively, as the context is clear. Given an arbitrary set of samples $S\in(\mathcal{X},\mathcal{Y})^{n}$ of size $n$ we write $\Phi_{\text{discounted}}(V)$ as follows.
$\Phi_{\text{discounted}}(V)$ $=\frac{|E(V,V)|-\sum_{u\in V}\|u-\mathcal{H}_{|S}\|_{0}}{|V|}$
$=\frac{|E(V,V)|-\sum_{u\in V}\mathop{min}_{h\in\mathcal{H}_{|S}}\|u-h\|_{0}}{|V|}$ $(\text{definition of }\|u-\mathcal{H}_{|S}\|_{0})$
$=\frac{n\cdot 2^{n-1}-\sum_{u\in V}\mathop{min}_{h\in\mathcal{H}_{|S}}\|u-h\|_{0}}{2^{n}}$ $(|E|=n\cdot 2^{n-1},|V|=2^{n})$
$=\frac{n}{2}-\frac{1}{2^{n}}\sum_{u\in V}\mathop{min}_{h\in\mathcal{H}_{|S}}\|u-h\|_{0}.$
Next, we switch notation from Hamming distance to inner product by using the following equality.
$\langle u,h\rangle:=\sum_{i=1}^{n}u(x_{i})\cdot h(x_{i})=\sum_{i=1}^{n}[u(x_{i})=h(x_{i})]-\sum_{i=1}^{n}[u(x_{i})\neq h(x_{i})]=n-2\cdot\|u-h\|_{0}.$ (2)
Then,
[…]
$=\frac{n}{2}-\frac{n}{2}-\frac{1}{2^{n}}\sum_{u\in V}\mathop{min}_{h\in\mathcal{H}_{|S}}-\frac{\langle u,h\rangle}{2}$ $(|V|=2^{n})$
$=\frac{1}{2}\cdot\frac{1}{2^{n}}\sum_{u\in V}\mathop{max}_{h\in\mathcal{H}_{|S}}\langle u,h\rangle$
$=\frac{1}{2}\cdot\mathop{\mathbb{E}}_{u\sim V}\left[\mathop{max}_{h\in\mathcal{H}_{|S}}\langle u,h\rangle\right].$
In the last step $u$ is sampled uniformly random from $V$. Notice that $u(x_{i})$ equals $+1$ or $-1$ with probability $1/2$ independently for any $(x_{i},y_{i})\in S$. Therefore,
$\Phi_{\text{discounted}}(V)=\frac{n}{2}\cdot\widehat{\Re}_{S}(\mathcal{H}_{|S})$
where ^$\widehat{\Re}_{S}(\mathcal{H}_{|S})$ is the empirical Rademacher complexity which is defined as:
^$\widehat{\Re}_{S}(\mathcal{H}_{|S}):=\frac{1}{n}\mathop{\mathbb{E}}_{\sigma\sim\{+1,-1\}^{n}}\left[\sup_{h\in\mathcal{H}_{|S}}\sum_{i=1}^{n}\sigma_{i}\cdot h(x_{i})\right].$
Finally, we invoke Lemma 10 to complete the proof:
$\Phi_{\text{discounted}}(V)=\frac{n}{2}\cdot\widehat{\Re}(\mathcal{H}_{|S})\leq 16\cdot\sqrt{n\cdot\operatorname{VC}(\mathcal{H})}.$
Appendix A Omitted proofs
A.1 Omitted proofs in Section 3
See 13
Proof.
Let $M_{i}=L_{\mathcal{D},\mathcal{H}}(h_{i-1})-d_{i}$. Then, as we observed above, the following conditional expectation evaluates to $\mathop{\mathbb{E}}[M_{i}\mid S_{i-1}]=L_{\mathcal{D}}(h_{i-1})-\mathop{\mathbb{E}}[d_{i}\mid S_{i-1}]=0.$ Then, random variables $M_{i}$ satisfies that $M_{i}\in[0,1]$ and $\mathop{\mathbb{E}}[M_{i}\mid M_{1},\dots M_{i-1}]\leq\epsilon(n)$ where $\epsilon(n)$ is the error rate of the transductive learner $A_{\text{Trans}}$.
Invoking Lemma 11 gives us
$\mathop{\mathbf{Pr}}\left[\sum_{i\in[k]}M_{i}\geq 2\cdot k\cdot\epsilon(n)\right]<\exp\left(-\frac{k\cdot\epsilon(n)}{3}\right).$
As we have $k=\frac{3\cdot\log(2/\delta)}{\epsilon}$, the claim follows. ∎
See 14
Proof.
Recall that we think of constructing our sequence of samples $S_{0},\dots,S_{k}$ “backwards” by first conditioning on the (unordered) set $S_{k}$, then iteratively peeling off one sample at a time — uniformly at random without replacement — to obtain $S_{k-1},S_{k-2},\ldots,S_{0}=S$. For $i=k,\ldots,1$, observe that $(x_{i},y_{i})$ is a uniformly random element from $S_{i}$, and that $|S_{i}|\geq n$. The transductive error assumption implies that $L_{S_{i}}^{tr}(A)\leq\epsilon(n)$ and therefore
$\mathop{\mathbb{E}}[d_{i}\mid S_{i}]\leq\epsilon.$
Since $d_{i}$ is conditionally independent of $d_{k},\ldots,d_{i+1}$ given $S_{i}$, it follows that
$\mathop{\mathbb{E}}[d_{i}\mid{d}_{k},\ldots,{d}_{i+1}]\leq\epsilon.$
As we have $d_{i}\in[0,1]$, invoking Lemma 11 once again gives us
$\mathop{\mathbf{Pr}}\left[\sum_{i\in[k]}d_{i}\geq 2\cdot k\cdot\epsilon(n)\right]\leq\exp\left(-\frac{k\cdot\epsilon(n)}{3}\right).$
Since $k=\frac{3\cdot\log(2/\delta)}{\epsilon}$, the claim follows. ∎
A.2 Omitted proofs in Section 4
See 18
Proof.
Until the end of the proof we condition on samples $U$. Denote $\Delta_{i}=d_{i}-\widetilde{d}_{i}$ and $\overline{\Delta}_{i}=\sum_{j=i}^{k}\Delta_{i}$. By definition of $d_{i}$ and $\widetilde{d}_{i}$, we have
$\Delta_{i}=[h^{*}_{S_{i}}(x_{i})\neq y_{i}]-[h^{*}_{\mathcal{D}}(x_{i})\neq y_{i}].$
Since $h^{*}_{S_{i}}$ is an optimal hypothesis in $\mathcal{H}$ for the empirical distribution $S_{i}$, the predictor $h^{*}_{\mathcal{D}}$ is in $\mathcal{H}$, and $(x_{i},y_{i})$ is uniformly distributed in $S_{i}$, it follows that
$\mathop{\mathbb{E}}[\Delta_{i}\mid S_{i}]\leq 0.$
Since $\Delta_{i}$ is conditionally independent of $\Delta_{i+1},\ldots,\Delta_{k}$ given $S_{i}$, it follows that
$\mathop{\mathbb{E}}[\Delta_{i}\mid\Delta_{i+1},\ldots,\Delta_{k}]\leq 0,$
or equivalently
$\mathop{\mathbb{E}}[\overline{\Delta}_{i}\mid\overline{\Delta}_{i+1}]\leq\overline{\Delta}_{i+1}.$
The sequence $\overline{\Delta}_{i}$ therefore forms a backwards super-martingale.
Invoking Azuma’s inequality together with the fact that $\Delta_{i}\in[-1,1]$, we obtain
$\mathop{\mathbf{Pr}}\left[\overline{\Delta}_{1}>\sqrt{8k\log\left(\frac{3}{\delta}\right)}\right]\leq\delta/3,$
as claimed. ∎
Appendix B Selecting a Hypothesis Using A Validation Set Gives a Low Error Hypothesis With High Probability
Lemma 21.
Let $H:=\{h_{0},\dots h_{k-1}\}\subseteq\mathcal{X}^{\mathcal{Y}}$ be a collection of predictors satisfying
$\mathop{\mathbf{Pr}}\left[\frac{1}{k}\cdot\sum_{i=0}^{k-1}L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(h_{i})>\epsilon\right]\leq\delta$
for some $\epsilon,\delta\in[0,1]$, and let $S_{\text{val}}:=(\mathcal{X}\times\mathcal{Y})^{t}$ be a cross-validation set of size $t=\frac{\log(k/\delta)}{\epsilon^{2}}$, sampled i.i.d. from the distribution $\mathcal{D}$. Then, for the predictor $\widehat{h}:=\mathop{argmin}_{h_{i}\in H}L_{S_{\text{val}},\mathcal{H}}^{\text{Ag}}(h_{i})$, which minimizes the agnostic loss on the cross-validation set, we have
$\mathop{\mathbf{Pr}}\left[L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(\widehat{h})>3\epsilon\right]\leq 2\cdot\delta.$
Proof.
Let $S_{\text{val}}:={(x_{1},y_{1}),\dots,(x_{t},y_{t})}$. Let $h^{*}\in H$ be the predictor with the minimum distributional loss among predictors in $H$. By assumption, with probability $1-\delta$, the average distributional loss of predictors $h\in H$ is at most $\epsilon$. From now on, we condition on this event and assume that the average agnostic distributional error is at most $\epsilon$. Therefore, optimal $h^{*}$ guarantees that $L_{D,\mathcal{H}}^{\text{Ag}}(h^{*})\leq\epsilon$. For $h^{*}$, define $X_{i}:=\ell(h^{*}(x_{i}),y_{i})$ for $i\in[t]$.
Using Hoeffding’s inequality and the fact that $X_{i}\leq 1$ for each $i\in[t]$, we observe that
$\mathop{\mathbf{Pr}}\left[L_{S,\mathcal{H}}^{\text{Ag}}(h^{*})\geq\epsilon+\sqrt{\frac{\log(1/\delta)}{2\cdot t}}\right]=\mathop{\mathbf{Pr}}\left[\frac{1}{t}\sum_{i=1}^{t}X_{i}\geq\epsilon+\sqrt{\frac{\log(1/\delta)}{2\cdot t}}\right]\leq\delta.$ (3)
On the other hand, let $h\in H$ be an arbitrary predictor with $L_{D,\mathcal{H}}^{\text{Ag}}(h)\geq\epsilon+2\cdot\sqrt{\frac{\log(k/\delta)}{2\cdot t}}$. Using Höeffding’s bound again, we observe that
$\mathop{\mathbf{Pr}}\left[L_{S,\mathcal{H}}^{\text{Ag}}(h)\leq\epsilon+\sqrt{\frac{\log(k/\delta)}{2\cdot t}}\right]\leq\frac{\delta}{k}.$
By the union bound, we can conclude that with probability at least $1-\delta$, there is no predictor $h\in H$ with $L_{S,\mathcal{H}}^{\text{Ag}}(h)\leq\epsilon+\sqrt{\frac{\log(k/\delta)}{2\cdot t}}$ and $L_{D,\mathcal{H}}^{\text{Ag}}(h)\geq\epsilon+2\cdot\sqrt{\frac{\log(k/\delta)}{2\cdot t}}$. By (3), we know that with probability $1-\delta$, there exists a predictor whose agnostic empirical loss is at most $\epsilon+\sqrt{\frac{\log(k/\delta)}{2\cdot t}}$. Therefore, we conclude that $\widehat{h}$, the empirically optimal predictor, attains at most $\epsilon+2\cdot\sqrt{\frac{\log(k/\delta)}{2\cdot t}}$ agnostic distributional error.
Finally, we release the condition on the average distributional error being at most $\epsilon$ and conclude that
$\mathop{\mathbf{Pr}}\left[L_{\mathcal{D},\mathcal{H}}^{\text{Ag}}(\widehat{h})\geq\epsilon+2\cdot\sqrt{\frac{\log(k/\delta)}{2\cdot t}}\right]\leq 3\cdot\delta.$
The proof is complete as we set $t=\frac{\log(k/\delta)}{\epsilon^{2}}$. ∎
Appendix C Connections to the results of Aden-Ali et al. [1]
In this section, we explore the connection between our work and that of Aden-Ali et al. [1]. Both papers present techniques for converting transductive learners to PAC learners in the realizable setting, but they make different assumptions about the properties of the underlying learners. Aden-Ali et al. [1] train multiple transductive learners by taking subsamples of the given sample set and aggregating their information to obtain a PAC learner. Their work explicitly defines this aggregation step for binary classification, partial concept classification, and bounded loss regression problems. In contrast, our approach involves training multiple learners by providing additional data to each and aggregating their outputs to form a PAC learner. Our aggregation method is more generic and applicable to any learning problem with bounded pseudometric loss functions.
Both our paper and the paper of Aden-Ali et al. [1] make assumptions on the worst case performance of a learner $A$ based on the size of the input sample. Intuitively, this paper assumes that when a learner is given larger samples, its expected error on those samples is non-increasing. That is,
$\epsilon_{\text{Trans}}(n)\leq\epsilon_{\text{Trans}}(n-1).\$ (A1)
In contrast, the authors of Aden-Ali et al. [1] assume that given smaller samples, the expected error of a learner does not increase more than linearly. More formally,
$n\epsilon_{\text{Trans}}(n)\geq(n-1)\epsilon_{\text{Trans}}(n-1).$ (A2)
As you may notice, these two assumptions are incomparable, or equivalently, one does not imply the other.
In this section, we will first show that A1 is without loss of generality true for transductive learners. We then show that assuming both (A1) and (A2) hold for transductive learners, the two results are equivalent up to constant factors.
C.1 The two assumptions in Section 2.1 are true without loss of generality
Lemma 22.
If a learner $\widetilde{A}$ achieves (agnostic) transductive error at most $\widetilde{\epsilon}(n)$, then there exists a randomized learner $A$ with transductive error rate of at most $\epsilon(n)$ such that $\epsilon(n+1)\leq\widetilde{\epsilon}(n)$.
Proof.
For any sample $S\in(\mathcal{X},\mathcal{Y})^{n+1}$ provided for transductive learning task. When a transductive learner $A$ is tested against data point $(x_{i},y_{i})$, $A$ subsamples a set $S^{\prime}\subseteq S_{-i}$ of size $n-1$ uniformly random and returns the label $A^{\prime}(S^{\prime})(x_{i})$. We can upper-bound transductive error of $A$ as follows.
$L_{S}^{\text{Trans}}(A)$ $=\mathop{\mathbb{E}}_{i\in[n]}\left[\mathop{\mathbb{E}}_{j\in[n]\setminus\{i\}}\left[\ell(A^{\prime}(S_{-i,-j})(x_{i}),y_{i})\right]\right]$
$=\mathop{\mathbb{E}}_{i\in[n]}[L_{S_{-i}}^{\text{Trans}}(A^{\prime})]$
$\leq\mathop{\mathbb{E}}_{i\in[n]}[\widetilde{\epsilon}(n)]$
$=\widetilde{\epsilon}(n).\qed$
Notice that the proof holds even for agnostic transductive learning problem when we replace the transductive error function with the agnostic transductive error function.
Moreover, for transductive learners with VC dimension $d$, we know that $\epsilon(n)=O\left(\frac{d}{n}\right)$ and $\epsilon_{\text{Ag}}(n)=O\left(\sqrt{\frac{d}{n}}\right)$. Therefore, Assumption A1 holds for these learners as well.
Let us say that we have a learner $A$ which is not symmetric and attains transductive error guarantee $L_{S}^{\text{Trans}}(A)\leq\epsilon_{\text{Trans}}$. We note that we can construct a learner $A^{\prime}$, where $A^{\prime}(S)=A(\pi(S))$ where $\pi$ is a permutation induced by sorting according to a consistent total ordering on $\mathcal{X}$. Note that because $A(S)\leq\epsilon_{\text{Trans}}$ regardless of the order of $S$, we see that $A(\pi(S))\leq\epsilon_{\text{Trans}}$ as well. In addition, we can see that $A^{\prime}$ is symmetric over input samples, as $\pi$ sorts the sample consistently. Therefore, we can conclude that if a learner $A$ exists attaining transductive error rate $\epsilon_{\text{Trans}}$, then a symmetric learner $A^{\prime}$ exists which also attains transductive error $\epsilon_{\text{Trans}}$.
C.2 Under Both of The Above Assumptions, Our Result is Equivalent to Aden-Ali et al. [1] Up To Constant Factors
We will show that if both assumptions (A2) and (A1) holds then Theorem 3 is equivalent up to constant factors to the main result of Aden-Ali et al. [1]. In particular, Aden-Ali et al. [1] demonstrated that given a learner $A$ which attains transductive error at most $\epsilon(n)$ on samples of size $n$, by training on $\frac{3n}{4}$ subsamples of that data, they demonstrated that under (A2), there are predictors $h_{1},\dots h_{3n/4}$ such that
$\mathop{\mathbf{Pr}}\left[\frac{4}{3n}\cdot\sum_{i=0}^{3n/4}L_{\mathcal{D}}(h_{i})\geq O\left(\epsilon(n)+\frac{\log(2/\delta)}{n}\right)\right]\leq\delta$ (4)
On the other hand, Lemma 12 states a similar probabilistic bound:
$\mathop{\mathbf{Pr}}\left[\frac{1}{k}\sum_{i=1}^{k}L_{\mathcal{D}}(h_{i})\geq O\left(\epsilon(n)+\frac{\log(2/\delta)}{k}\right)\right]\leq\delta.$
Note that two claim are equivalent to each other when the additional data size $k$ is set to $3n$ in Lemma 12.